# Accelerating shape optimization by deep neural networks with on-the-fly determined architecture

### v1

@@ Abstract
In component shape optimization, the component properties are often evaluated by computationally expensive simulations. Such optimization becomes unfeasible when it is focused on a global search requiring thousands of simulations to be evaluated. Here, we present a viable global shape optimization methodology based on multi-objective evolutionary algorithms accelerated by deep neural networks (DNNs). Our methodology alternates between evaluating simulations and utilizing the generated data to train DNNs with various architectures. When a suitable DNN architecture is identified, the DNN replaces the simulation in the rest of the global search. Our methodology was tested on five ZDT benchmark functions, showing itself at the level of and sometimes more flexible than other state-of-the-art acceleration approaches. Then, it was applied to a real-life optimization problem, namely the shape optimization of a single-phase ejector. Compared with a non-accelerated methodology, ours was able to save weeks of CPU time in solving this problem. To experimentally confirm the performance of the optimized ejector shapes, four of them were 3D printed and tested on the lab scale confirming the predicted performance. This suggests that our methodology could be used for acceleration of other real-life shape optimization problems.
@@ 1 Introduction
Our methodology is first benchmarked using the ZDT functions proposed by zitzler2000 for testing multi-objective optimization algorithms. Its performance is compared with other similar state-of-the-art SAOAs [muller2017, morita2022, mastrippolito2021]. Next, the methodology is used to solve a real-life optimization problem; namely, the shape optimization of a single-phase ejector. The motivation for this problem choice was twofold. First, it represents a relatively simple task for CFD simulations. Second, we have access to a specifically tailored experimental setup that was eventually used to validate the optimization results. Furthermore, the performance of the ejector is known to strongly depend on its geometry, and the effects of alterations in the geometry of its various parts on its performance have been investigated in recent decades [havelka1997, utomo2009, daniels2019]. In our case, the shape optimization focused on the converging part and diffuser because these parts can be easily 3D printed and their optimized performance experimentally validated.
The paper proceeds as follows. First, multi-objective optimization is introduced in general, which is followed by a description of multi-objective evolutionary algorithms. Next, we provide details on our surrogate-assisted optimization algorithm. The algorithm is based on an adaptive connection between DNNs and CFD and is hereafter abbreviated as CFDNNetAdapt. In the subsequent part, the benchmarking using ZDT functions is introduced. Lastly, the real-life optimization problem is described in detail, and the optimization results are discussed. Details on the experimental setup and the used CFD model are given in the appendix.
@@ 2.1 Multi-objective optimization problem
In case the objectives are anti-correlated, i.e., improving performance in one objective may worsen performance in the others, there does not exist a single optimal solution. Instead, a set of trade-off solutions is sought. This set is commonly referred to as the Pareto-optimal set ($\mathcal{P}^{*}$). Its counterpart with trade-off objective values is the Pareto-optimal front ($\mathcal{O}^{*}$) [pareto1896].
@@ 2.2 Multi-objective evolutionary algorithms
To solve a multi-objective optimization problem, we opted for using the multi-objective evolutionary algorithms (MOEAs) [emmerich2018]. Of the many MOEA variants available, we used here the NSGA-II algorithm by deb2002. The algorithm works with a population of individua that is updated each iteration. Each individuum is a tuple $(\bm{p},\bm{f}_{\mathrm{cost}}(\bm{p}))$. The population size ($n_{\mathrm{pop}}$) is preset and the first population consists of individua with randomly generated vectors $\bm{p}$ within a range defined by $\bm{p}_{\mathrm{min}}$ and $\bm{p}_{\mathrm{max}}$.
In each iteration, random pairs of individua are chosen and mixed using partially random operators such as crossover and mutation, generating offspring individua with different $\bm{p}$ vectors. The offspring individua are included in the population, and all the individua are sorted. In NSGA-II, the sorting is based on the Pareto-dominance principle. Lastly, the population size is reduced to $n_{\mathrm{pop}}$ by truncating the least fit individua identified by the sorting [deb2002]. After this truncation, a new iteration starts. Note that each iteration is also referred to as a generation. The algorithm terminates after a preset number of generations ($n_{\mathrm{gen}}$).
@@ 2.3 Baseline optimization framework
Figure 1: Schematic of the baseline optimization framework. Regions operated by either OpenFOAM or Python are distinguished. The vectors $\bm{p}$ and $\bm{o}$ contain the parameters and objectives, respectively. Furthermore, $\bm{f}_{\mathrm{cost}}$ is the cost function, $\mathcal{P}^{*}$ the Pareto-optimal set and $\mathcal{O}^{*}$ the Pareto-optimal front.
The core of the framework is implemented in Python. NSGA-II is included through an open-source Python library for multi-objective evolutionary optimization Platypus [platypus]. This library is combined with a custom code for the $f_{\mathrm{cost}}$ evaluation. For our real-life application, the $f_{\mathrm{cost}}$ evaluation comprises an automated preparation of a CFD model in OpenFOAM [OpenFOAM2007]. The description of the CFD model construction for our real-life application (optimization of an ejector pump) is given in B.
The framework is able to evaluate individua in each generation in parallel. In addition, the CFD models can also be evaluated in parallel. This allows for an effective run on high-performance computing architectures.
@@ 3.1 Multilayer perceptrons
Regarding implementation, everything related to MLPs, e.g., construction or subsequent supervised learning, was implemented using an open-source Python library pyrenn [pyrenn].
@@
hyperparameter description values used in Section 4
values used in Section 5
$n_{\mathrm{hlrs}}$ number of hidden layers 33
33
$n_{\mathrm{dnn}}$ number of MLPs generated 44
in each iteration $10$
$r_{n}$ half width of the interval 44
for choosing hidden layer size 44
$n_{\mathrm{max}}$ maximal size of MLP layer 2020
2020
$n_{\mathrm{min}}$ minimal size of MLP layer 22
22
$\bar{n}^{0}_{i}$ initial value for the average of $11$ for any $i$
the $i$-th hidden layer size $11$ for any $i$
$n_{\mathrm{smp}}$ number of data points $1000$
generated in one iteration $600$
$n_{\mathrm{pop}}$ population size $100$
for NSGA-II $500$
$n_{\mathrm{gen}}$ number of generations $250$
for NSGA-II $90$
$n_{\mathrm{ver}}$ number of points chosen $16$
for verification 88
$\epsilon$ required tolerance $10^{-4}$
$0.05$
$k_{\mathrm{max}}$ maximum number of $25$
iterations $12$
@@ 3.2 CFDNNetAdapt
where $n_{\mathrm{max}}$ and $n_{\mathrm{min}}$ are hyperparameters defining the maximal and minimal allowed layer size, respectively. Note that the interval half width is kept the same for all the hidden layers, but the average value can be different. In the first iteration, the averages are preset (hyperparameters $\bar{n}^{0}_{i}$, $i=1,\,\dots,\,n_{\mathrm{hlrs}}$). In subsequent iterations, the averages are inherited from the best MLP found in the previous iteration.
@@
All data stored in the data bank are processed together. First, non-dominated points are identified and saved as an intermediate Pareto-optimal set, $\mathcal{P}^{*}_{\mathrm{CFD}}$. Next, all data points are randomly divided into three groups for supervised learning of MLPs. Specifically, $75\,\%$ are labeled as training data, $15\,\%$ as validation data, and $10\,\%$ as testing data.
@@
Each MLP is trained on the training data using the Levenberg-Marquardt algorithm [levenberg1944, marquardt1963]. The validation data are used to track MLP performance during training, and if the performance worsens significantly, training is stopped to prevent overfitting [goodfellow2016]. After the training is completed, testing data are used to evaluate the performance of each MLP, and the results are saved to allow for during-run or after-run checks of the algorithm’s progress. Inclusion of the performance check based on the testing data is motivated by an effort to allow the user to identify the potential overfitting of adaptively selected MLP architectures to the validation data.
@@
Each trained MLP is used as a surrogate in an optimization run by NSGA-II. Each optimization runs with a predefined population size (hyperparameter $n_{\mathrm{pop}}$) and number of generations (hyperparameter $n_{\mathrm{gen}}$). Hence, each MLP predicts its own Pareto-optimal set ($\mathcal{P}^{*}_{\mathrm{MLP}}$) and front ($\mathcal{O}^{*}_{\mathrm{MLP}}$).
@@
The quality of each MLP-predicted Pareto-optimal set ($\mathcal{P}^{*}_{\mathrm{MLP}}$) is estimated by comparing it with the intermediate Pareto-optimal set identified from the data points ($\mathcal{P}^{*}_{\mathrm{CFD}}$). The comparison is based on the evaluation of an inverted generational distance (IGD) defined as
$\text{IGD}(\mathcal{P}^{*}_{\mathrm{CFD}},\,\mathcal{P}^{*}_{\mathrm{MLP}})=\frac{1}{\|\mathcal{P}^{*}_{\mathrm{CFD}}\|}\displaystyle\sum_{\bm{x}\in\mathcal{P}^{*}_{\mathrm{CFD}}}\min_{\bm{y}\in\mathcal{P}^{*}_{\mathrm{MLP}}}\left\lVert\bm{x},\,\bm{y}\right\rVert\,$, (4)
where the lower the IGD, the more similar the sets are. As the most suitable Pareto-optimal set, the one with the lowest IGD is chosen. This choice, which may be seen as method under-relaxation, originates from the decreasing reliability of sets with a higher IGD. When a poorly trained MLP is used in an optimization, it tends to exploit gaps in its training and come up with solutions that are far off from what it was trained on, thus potentially having a large prediction error. Therefore, the Pareto-optimal set with the lowest IGD is assumed to contain the lowest number of such exploited solutions.
@@ 4 Benchmarking of CFDNNetAdapt
For the verification of CFDNNetAdapt, five benchmark optimization problems suggested by zitzler2000 were chosen. The problems are referred to as ZDT1, ZDT2, ZDT3, ZDT4, and ZDT6. All of these operate with real parameters and two objectives. The test function ZDT5 was not included as its parameters are decoded as bitstrings. The complete specification of the test functions chosen is given in Table 2. Optimal solutions to each problem are illustrated in black in Figure 3.
@@
The performance of CFDNNetAdapt was compared with two widely used algorithms: (i) NSGA-II without any surrogate model [deb2002], and (ii) state-of-the-art surrogate-assisted optimization algorithm SOCEMO [muller2017]. In addition, a variant of CFDNNetAdapt with non-DNN surrogate model was constructed and tested to investigate the suitability of DNN usage. A more detailed description of each method settings is provided in the following paragraphs, and the results are depicted in Figures 3 and 4.

### v2

@@ Abstract
In component shape optimization, the component properties are often evaluated by computationally expensive simulations. Such optimization becomes unfeasible when it is focused on a global search requiring thousands of simulations to be evaluated. Here, we present a viable global shape optimization methodology based on multi-objective evolutionary algorithms accelerated by deep neural networks (DNNs). Our methodology alternates between evaluating simulations and utilizing the generated data to train DNNs with various architectures. When a suitable DNN architecture is identified, the DNN replaces the simulation in the rest of the global search. Our methodology was tested on a number of standardized benchmark functions, showing itself at the level of and sometimes more flexible than other state-of-the-art acceleration approaches. Then, it was applied to a real-life optimization problem, namely the shape optimization of a single-phase ejector. Compared with a non-accelerated methodology, ours was able to save weeks of CPU time in solving this problem. Additionally, four optimized ejector shapes were 3D printed and tested experimentally to show their real-life performance.
@@ 1 Introduction
Our methodology is first benchmarked using the ZDT functions of zitzler2000 and LZ functions by li2009 proposed for testing multi-objective optimization algorithms. Its performance is compared with other similar state-of-the-art SAOAs [muller2017, morita2022, mastrippolito2021]. Next, the methodology is used to solve a real-life optimization problem; namely, the shape optimization of a single-phase ejector. The motivation for this problem choice was twofold. First, it represents a relatively simple task for CFD simulations. Second, we have access to a specifically tailored experimental setup that was eventually used to validate the optimization results. Furthermore, the performance of the ejector is known to strongly depend on its geometry, and the effects of alterations in the geometry of its various parts on its performance have been investigated in recent decades [havelka1997, utomo2009, daniels2019]. In our case, the shape optimization focused on the converging part and diffuser because these parts can be easily 3D printed and their optimized performance experimentally validated.
The paper proceeds as follows. In Section 2, multi-objective optimization is introduced in general and a description of multi-objective evolutionary algorithms is given. In Section 3, we provide details on our surrogate-assisted optimization algorithm. The algorithm is based on an adaptive connection between DNNs and CFD and is hereafter abbreviated as CFDNNetAdapt. In Section 4, the benchmarking using ZDT functions is introduced, with further benchmarks based on LZ functions provided in Supplementary material. Lastly, in Section 5, the real-life optimization problem is described in detail, and the optimization results are discussed. Details on the experimental setup and the used CFD model are given in the appendix.
@@ 2.1 Multi-objective optimization problem
In case the objectives are anti-correlated, i.e., improving performance in one objective may worsen performance in the others, there does not exist a single optimal solution. Instead, a set of trade-off solutions is sought. This set is commonly referred to as the Pareto-optimal set ($\mathcal{P}^{*}$). Its counterpart with trade-off objective values is the Pareto-optimal front ($\mathcal{O}^{*}$[pareto1896].
@@ 2.2 Multi-objective evolutionary algorithms
To solve a multi-objective optimization problem, we opted for using the multi-objective evolutionary algorithms (MOEAs) [emmerich2018]. Of the many MOEA variants available, we used here the NSGA-II algorithm by deb2002. The algorithm works with a population of individuals that is updated each iteration. Each individual is a tuple $(\bm{p},\bm{f}_{\mathrm{cost}}(\bm{p}))$. The population size ($n_{\mathrm{pop}}$) is preset and the first population consists of individuals with randomly generated vectors $\bm{p}$ within a range defined by $\bm{p}_{\mathrm{min}}$ and $\bm{p}_{\mathrm{max}}$.
In each iteration, random pairs of individuals are chosen and mixed using partially random operators such as crossover and mutation, generating offspring individuals with different $\bm{p}$ vectors. The offspring individuals are included in the population, and all the individuals are sorted. In NSGA-II, the sorting is based on the Pareto-dominance principle. Lastly, the population size is reduced to $n_{\mathrm{pop}}$ by truncating the least fit individuals identified by the sorting [deb2002]. After this truncation, a new iteration starts. Note that each iteration is also referred to as a generation. The algorithm terminates after a preset number of generations ($n_{\mathrm{gen}}$).
@@ 2.3 Baseline optimization framework
Figure 1: Schematic of the baseline optimization framework. The vectors $\bm{p}$ and $\bm{o}$ contain the parameters and objectives, respectively. Furthermore, $\bm{f}_{\mathrm{cost}}$ is the cost function, $\mathcal{P}^{*}$ the Pareto-optimal set and $\mathcal{O}^{*}$ the Pareto-optimal front.
@@ 3.1 Multilayer perceptrons
hyperparameter description
$n_{\mathrm{hlrs}}$ number of hidden layers
$n_{\mathrm{dnn}}$ number of MLPs generated in each iteration
$r_{n}$ half width of the interval for choosing hidden layer size
$n_{\mathrm{max}}$ maximal size of MLP layer
$n_{\mathrm{min}}$ minimal size of MLP layer
$\bar{n}^{0}_{i}$ initial value for the average of the $i$-th hidden layer size
$n_{\mathrm{smp}}$ number of data points generated in one iteration
$n_{\mathrm{pop}}$ population size for NSGA-II
$n_{\mathrm{gen}}$ number of generations for NSGA-II
$n_{\mathrm{ver}}$ number of points chosen for verification
$\epsilon$ required tolerance
$k_{\mathrm{max}}$ maximum number of iterations
@@ 3.2 CFDNNetAdapt
where $n_{\mathrm{max}}$ and $n_{\mathrm{min}}$ are hyperparameters defining the maximal and minimal allowed layer size, respectively. Note that the interval half-width is kept the same for all the hidden layers, but the average value can be different. In the first iteration, the averages are preset (hyperparameters $\bar{n}^{0}_{i}$, $i=1,\,\dots,\,n_{\mathrm{hlrs}}$). In subsequent iterations, the averages are inherited from the best MLP found in the previous iteration.
@@
All data stored in the data bank are processed together. First, non-dominated points are identified and saved as an intermediate Pareto-optimal set, $\mathcal{P}^{*}_{\mathrm{CFD}}$. Next, all data points are sorted into three groups for supervised learning of MLPs. Specifically, the data points are randomly shuffled and the first $75\,\%$ are labeled as training data, the next $15\,\%$ as validation data, and the last $10\,\%$ as testing data.
@@
Each MLP is trained on the training data using the Levenberg-Marquardt algorithm [levenberg1944, marquardt1963]. To prevent overfitting, the validation data are used to monitor the MLP performance during training [goodfellow2016] and the training termination is decided based on non-successful training iterations. In particular, after each training iteration, the MLP is used to evaluate the validation data points, and the mean error between the predicted and true values is computed. If the error is greater than the error from the previous iteration, the MLP receives one strike. If the total number of strikes becomes greater than a threshold value (value of $10\cdot n_{\mathrm{hlrs}}$ was used in this work), the training is terminated.
After the training is completed, testing data are used to evaluate the performance of each MLP, and the results are saved to allow for during-run or after-run checks of the algorithm progress. Inclusion of the performance check based on the testing data is motivated by an effort to allow the user to identify the potential overfitting of adaptively selected MLP architectures to the validation data.
@@
With each trained MLP, a separate optimization is run where the MLP acts as a surrogate model in cost function evaluation. The optimizations are run by NSGA-II with a predefined population size (hyperparameter $n_{\mathrm{pop}}$) and number of generations (hyperparameter $n_{\mathrm{gen}}$). Hence, each MLP predicts its own Pareto-optimal set ($\mathcal{P}^{*}_{\mathrm{MLP}}$) and front ($\mathcal{O}^{*}_{\mathrm{MLP}}$).
@@
From the $n_{\mathrm{dnn}}$ trained MLPs, the best one has to be selected. This selection is based on the similarity of the MLP-predicted Pareto optimal sets ($\mathcal{P}^{*}_{\mathrm{MLP}}$) to the intermediate Pareto-optimal set ($\mathcal{P}^{*}_{\mathrm{CFD}}$) identified from the data points. The MLP that predicted the most similar Pareto-optimal set is chosen as the best one.
The similarity of the Pareto-optimal sets is evaluated using the inverted generational distance (IGD) defined in agreement with [li2009] as
$\text{IGD}(\mathcal{P}^{*}_{\mathrm{CFD}},\,\mathcal{P}^{*}_{\mathrm{MLP}})=\frac{1}{\|\mathcal{P}^{*}_{\mathrm{CFD}}\|}\displaystyle\sum_{\bm{x}\in\mathcal{P}^{*}_{\mathrm{CFD}}}\min_{\bm{y}\in\mathcal{P}^{*}_{\mathrm{MLP}}}\left\lVert\bm{x},\,\bm{y}\right\rVert\,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}.}$ (4)
The Pareto-optimal set $\mathcal{P}^{*}_{\mathrm{MLP}}$ with the lowest IGD value is perceived as the most similar, and the corresponding MLP is chosen as the best MLP.
This choice, which may be seen as a method of under-relaxation, originates from the decreasing reliability of sets with a higher IGD. When a poorly trained MLP is used in an optimization, it tends to exploit gaps in its training and come up with solutions that are far off from what it was trained on, thus potentially having a large prediction error. Therefore, the Pareto-optimal set with the lowest IGD is assumed to contain the lowest number of such exploited solutions.
@@
Note on stochastic nature of CFDNNetAdapt
First, CFDNNetAdapt uses a MOEA (in our case NSGA-II) as a data generator. Thus, it inherently includes stochastic operations. Second, beyond the standard sources of randomness in MOEAs, CFDNNetAdapt also introduces randomness through variations in the sizes of the hidden layers of the MLPs used as surrogates. The random selection of hidden-layer sizes, followed by evaluating each MLP performance via IGD is motivated by the need to identify MLPs that have good generalization capabilities and are not prone to overfitting. However, this procedure adds an additional layer of stochasticity to the behavior of CFDNNetAdapt. To illustrate the effects of this randomness, further details and a statistical analysis of the variability in MLP architectures across all optimization problems considered in this work are provided in D.
3.3 Implementation details
To conclude the description of the CFDNNetAdapt algorithm, we briefly describe implementations of (i) the baseline optimization framework, and (ii) CFDNNetAdapt. In addition, CFDNNetAdapt is an open-source project, and all source codes and configuration files necessary to reproduce the computations presented in this study are available from https://github.com/techMathGroup/CFDNNetAdapt.
Specifically, both the baseline and the surrogate-aided frameworks are implemented in Python and prepared to run on high-performance computing architectures. The MOEAs are included through an open-source Python library for multi-objective evolutionary optimization Platypus [platypus]. Everything related to MLPs is implemented using an open-source Python library pyrenn [pyrenn]. These libraries are combined with a custom code for the evaluation of $f_{\mathrm{cost}}$. Moreover, for the purposes of the real-life application, the $f_{\mathrm{cost}}$ evaluation comprises an automated preparation of a CFD model in OpenFOAM [OpenFOAM2007]. The description of the CFD model construction for our real-life application (optimization of an ejector pump) is given in B.
@@ 4 Benchmarking of CFDNNetAdapt
For the verification of CFDNNetAdapt, several benchmark optimization problems were chosen including the ZDT functions suggested by zitzler2000 and LZ functions proposed by li2009. In this section, the results of verification tests using the ZDT functions are presented. The results obtained using the LZ functions are similar in nature to the results obtained using ZDT functions. Therefore, the verification results from the LZ functions are presented in the Supplementary material.
For the verification presented here, we used five of the ZDT functions which are referred to as ZDT1, ZDT2, ZDT3, ZDT4, and ZDT6. All of these operate with real parameters and two objectives. The test function ZDT5 was not included as its parameters are decoded as bitstrings. The complete specification of the test functions chosen is given in Table 2. Optimal solutions to each problem are illustrated in black in Figure 3.
@@
hyperpar. value hyperpar. value hyperpar. value
$n_{\mathrm{hlrs}}$ 33 $n_{\mathrm{min}}$ 22 $n_{\mathrm{gen}}$ $250$
$n_{\mathrm{dnn}}$ 44 $n_{\mathrm{max}}$ 2020 $n_{\mathrm{ver}}$ $16$
$\{\bar{n}^{0}_{1},\,\bar{n}^{0}_{2},\,\bar{n}^{0}_{3}\}$ {11, 11, $\{11$,\,11,\,11\} $n_{\mathrm{smp}}$ $1,000$ $k_{\mathrm{max}}$ $100$
$r_{n}$ 44 $n_{\mathrm{pop}}$ $100$ $\epsilon$ $10^{-6}$
Table 3: Values of CFDNNetAdapt hyperparameters used for the ZDT functions. Individual hyperparameters are explained in Table 1.
4.1 Compared algorithms and algorithm settings
@@
The performance of CFDNNetAdapt was compared with two widely used algorithms (i) NSGA-II without any surrogate model [deb2002], and (ii) state-of-the-art surrogate-assisted optimization algorithm SOCEMO [muller2017]. In addition, a variant of CFDNNetAdapt with non-DNN surrogate model was constructed to investigate the suitability of DNN usage. As the non-DNN surrogate model, we have chosen models based on Gaussian process regression (GPR) [jones1998]. The choice of GPR was motivated by its frequent use (mainly as Kriging models) in SAOAs, as described in, e.g. [morita2022, mastrippolito2021]. Hereafter, this variant will be referred to as CFDGPRAdapt. A more detailed description of each method settings is provided in the following paragraphs, and the results are depicted in Figures 3– 5.
Similarly to deb2002, NSGA-II was run with population size $n_{\mathrm{pop}}=250$ and number of generations $n_{\mathrm{gen}}=100$. All evaluated data points (tuples ($\bm{p}$$\bm{o}$)) generated by NSGA-II ($\sim\!\!25,000$) were used to feed the other SAOAs. The SAOAs training sets were gradually extended as follows: the first batch comprised $10$ samples, the second $30$, then $100;300;500;$ and $1,000;2,000;\dots;25,000$ samples.
The setting of CFDNNetAdapt hyperparameters is given in Table 3 and was kept the same for all the ZDT functions to have the comparison fair. To set the hyperparameters defining the MLP size ($n_{\mathrm{hlrs}}$, $n_{\mathrm{max}}$), a few MLPs with small to large architectures (increasing both $n_{\mathrm{hlrs}}$ and $n_{\mathrm{max}}$) were trained using from $200$ to $1,000$ NSGA-II samples generated for ZDT1. After the training, a mean error of the MLPs was evaluated on unseen NSGA-II samples. For $n_{\mathrm{hlrs}}\geq 3$ and $n_{\mathrm{max}}\geq 20$, the error stagnated at $\delta_{\mathrm{stg}}\sim 10^{-6}$. Based on this testing, the $n_{\mathrm{hlrs}}$ and $n_{\mathrm{max}}$ were set as the limit values. Also, the $\epsilon$ could be set as $\epsilon\sim\delta_{\mathrm{stg}}\approx 10^{-6}$ and $n_{\mathrm{ver}}$ was estimated to provide a trade-off between accuracy and computational costs.
Next, the hyperparameter $n_{\mathrm{min}}$ was set as the smallest possible value, i.e., $n_{\mathrm{min}}=2$, and the values of $\{\bar{n}_{1},\,\bar{n}_{2},\,\bar{n}_{3}\}$ and $r_{n}$ were computed as
$\bar{n}^{0}_{i}=\frac{1}{2}\left(n_{\mathrm{max}}+n_{\mathrm{min}}\right),\quad r_{n}=\frac{1}{4}\left(n_{\mathrm{max}}-n_{\mathrm{min}}\right),\,$ (6)
Lastly, the values of $n_{\mathrm{pop}}$ and $n_{\mathrm{gen}}$ were set the same as for pure NSGA-II. Based on the total number of NSGA-II samples ($\sim 25,000$), the $n_{\mathrm{smp}}$ was chosen as $1,000$ and the expected number of iterations is then $25$. The maximum number of iterations $k_{\mathrm{max}}$ was chosen as four times the expected number.
Next, the surrogates used by SOCEMO are based on radial basis functions (RBFs) and were originally used for problems with up to $500$ data points. As we work with up to $25,000$ points, we implemented a probabilistic selection of inducing points based on [uhrenholt2021] to effectively lower the problem dimension.
Finally, the setting of CFDGPRAdapt was kept as close as possible to the setting of CFDNNetAdapt presented in Table 3. Only the settings used for MLPs could not be used and new settings for GPRs was added. Specifically, the sparse variant of the GPR with up to $50$ inducing points was used. In each iteration, four GPR models with different kernels (squared exponential, Matern52, Matern32, and Matern12) were tested. The implementation was done using a Python library gpflow [gpflow2017].
4.2 Estimates of Pareto-optimal solutions